Item 17 Part A What capacitor in series with a 100 Ω resistor and a 26 mH inductor will give a resonance frequency of 1080 Hz ? Express your answer using two significant figures. HF Submit My Answers Give UpWhat capacitor in series with a 100 ? resistor and a 26 mH inductor will give a resonance frequency of 1080 Hz ? Express your answer using two significant figures.
Answer #1

given that

L = 26 mH = 26*10^(-3) H

resonance frequency f = 1080 Hz

we know that

resonance frequency f = 1/2*pi*sqrt(L*C)

C = 1 / 4*pi^2*f^2*L

C = 1 / 4*3.14^2*1080^2*26*10^(-3)

C = 0.836*10^(-6) F

C = 0.836 uF